# Scaling Limit of a Stochastic Clustering Model on $\mathds{R}$

### v1

@@ 2.3 Time Reversal of Gap Sequence Model
Recall that in the forward dynamics, $1/4$ of the points are lost due to merging. In the time-reversal, $1/3$ of points should thus split into two new points. Indeed, in the forward process, the inter-duration between points which merge with their left neighbor is given by the sum of two independent $\operatorname{Geom}(1/2)$ random variables. Thus, in the time-reversed process, points “un-merge” with inter-duration $\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1$, where the geometrics are again independent. Thus, let $\rho^{(t)}$ be i.i.d. renewal processes with gap distribution
$R\sim\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1$
where the geometric random variables are independent and take the minimum value $1$. We have $\operatorname{\mathds{E}}R=2+2-1=3$. The requisite coupling between the merged indices and the point processes $\rho^{(t)}$ used in the un-merging is obvious and we omit it for the sake of brevity; see instead Figure 3.
@@
Here, this summation corresponds to the fact that the same points contribute to two different gaps after one iteration of the time-reversed dynamics. The sequence $\overleftarrow{\eta}^{(t+1)}$ is indexed so that the 0–indexed point is the sum of the left-most element of the 0–th tuple and the right-most element of the (1)(-1)–st tuple. For example, in Figure 3, the un-mergings correspond to the entire gap which gets deleted by the merging, but also half of each of the preceding and following gaps.
@@ 2.4 Main Results
We do not actually know the distribution of the Palm shift. The limit is not renewal. Thus, it is difficult to identify and is left as an open problem. The following corollary is a consequence of the proof technique of Theorem 2.1
@@
The time-reversed process $\left({3}/{8}\right)^{t}\eta^{(t)}$ is the stochastic dual of the gap sequence model, with respect to the inner product
$h(\Gamma^{(s)},\left({3}/{8}\right)^{t}\eta^{(t)}):=\left\langle\Gamma^{(s)},\left({3}/{8}\right)^{t}\eta^{(t)}\right\rangle=\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\Gamma^{(s)}_{i}\eta^{(t)}_{i}$.
@@
Let $\overrightarrow{G}^{(t)}$ be the number of points of $\overrightarrow{\Xi}^{(t)}$ which have merged with the 0–indexed point of $\Xi^{(0)}$. Then
@@
for $2\leqslant p<\infty$.
The following corollary follows by coupling the objects in the Proofs of Theorems 2.1 and 2.3; but it is not a direct consequence of the theorem statements themselves.
@@
The joint weak convergence holds:
$\left(\Theta\overrightarrow{\Xi}^{(t)},\left(\frac{3}{4}\right)^{t}\overrightarrow{G}^{(t)}\right)\Rightarrow\left(\Theta\overrightarrow{\Xi}^{(\infty)},\overrightarrow{G}^{(\infty)},\right).$
@@ 3 Proofs of Main Results
Recall that $\overleftarrow{\eta}^{(t)}$ is a sequence of weights, from which we will construct (after a re-scaling by $\left(\frac{3}{8}\right)^{t}$ the gap sequence of the limiting point process. Note that $\overleftarrow{\eta}^{(t)}$ is supported on a subset of $\{-2^{t},-2^{t}+1,\ldots,2^{t}\}$, so that
$|\overleftarrow{\eta}^{(t)}|:=\sum_{i=-\infty}^{\infty}{\overleftarrow{\eta}^{(t)}_{i}}$$\infty$.
Define for $t\geqslant 0$ the process
@@
We can express the evolution of $\overleftarrow{M}^{(t)}$ as follows:
$\overleftarrow{M}^{(t+1)}$ $=\left(\frac{3}{8}\right)^{t+1}\left(2|\overleftarrow{\eta}^{(t)}|+2\sum_{i\in\rho^{(t+1)}}\overleftarrow{\eta}^{(t)}_{i}\right)$
@@
$=\overleftarrow{M}^{(t)}+\frac{3}{4}\sum_{i\in\mathrm{supp}\overleftarrow{\eta}^{(t)}}\overleftarrow{X}^{(t)}_{i}\left(\mathds{1}_{i\in\rho^{(t+1)}}-\frac{1}{3}\right)=:\overleftarrow{M}^{(t)}+\overleftarrow{N}^{(t+1)}.$
Here, $\overleftarrow{X}^{(t)}:=\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}$. Notice that $\overleftarrow{N}^{(t+1)}$ has conditional mean 0. It follows that the process $\overleftarrow{M}^{(t)}$ is a unit-mean positive martingale; hence it converges to an a.s. limit $\overleftarrow{M}^{(\infty)}$.
Notice that since $\max_{i\in\mathds{Z}}\overleftarrow{\eta}^{(t)}_{i}\leqslant 2^{t}$, we also have that $\max_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}\leqslant\left(\frac{3}{4}\right)^{t}$. Equivalently, we have $\max_{i\in\mathds{Z}}X_{i}^{(t)}\leqslant\left(\frac{3}{4}\right)^{t}$. Thus,
$\sum_{i\in\mathds{Z}}\left(\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta^{(t)}_{i}}\right)^{2}\leqslant\left(\frac{3}{4}\right)^{t}\overleftarrow{M}^{(t)},$
@@
since $\overleftarrow{N}^{(t+1)}$ is conditionally 0 mean. Substituting, we have:
@@
$\leqslant\sum_{i\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\sum_{j\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\left(\frac{3}{4}\right)^{2t}\operatorname{Cov}(\mathds{1}_{i\in\rho^{(t+1)}},\mathds{1}_{j\in\rho^{(t+1)}})$
@@
Here, the convergence occurs using the fact that the renewal distribution of $\rho$ has an exponential tail and thus, $\operatorname{Cov}(\mathds{1}_{i\in\rho^{(t)}},\mathds{1}_{j\in\rho^{(t)}})$ decays exponentially in |ji||j-i|. It follows that
@@
Figure 4: Relationship between the weight sequences $\eta^{(t)}$ and the initial gap sequence $\Gamma^{(0)}$. The orange dashed line, for example, is given by $\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{-1}+\overleftarrow{\Gamma}^{(0)}_{0}+\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{1}$; the blue dashed line is given by $\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{0}+\overleftarrow{\Gamma}^{(0)}_{1}+\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{2}$.
Now, if the gap distribution of $\Xi_{0}$ has finite variance, we have the following (see Figure 4):
$\left\langle\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)},\overleftarrow{\Gamma}^{(0)}\right\rangle$ $:=\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}\overleftarrow{\Gamma}^{(0)}_{i}=\overleftarrow{M}^{(t)}+\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}(\overleftarrow{\Gamma}^{(0)}_{i}-1)=:\overleftarrow{M}^{(t)}+\overleftarrow{L}^{(t)}.$
Obviously, we have $\operatorname{\mathds{E}}\overleftarrow{L}^{(t)}=0$ for all $t$, since $\operatorname{\mathds{E}}\Gamma^{(0)}_{i}=1$ for all $i\in\mathds{Z}$. Noting that $(\overleftarrow{\Gamma}_{i}^{(0)})_{i\in\mathds{Z}}$ are i.i.d. with variance $\sigma^{2}$ gaps, we have
$\operatorname{\mathds{E}}\left[\left(\overleftarrow{L}^{(t)}\right)^{2}\ \biggl|\ \mathcal{F}^{(t)}\right]$ $=\sum_{i\in\mathds{Z}}\sum_{j\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta_{i}}^{(t)}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}_{j}^{(t)}\cdot\operatorname{Cov}(\overleftarrow{\Gamma}_{i}^{(0)},\overleftarrow{\Gamma}_{j}^{(0)})$

### v2

@@ 3.3 Time Reversal of Gap Sequence Model
Recall that in the forward dynamics, $1/4$ of the points are lost due to merging. In the time reversal, $1/3$ of the points should thus split into two new points. Indeed, in the forward process, the interval between points that merge with their left neighbors is given by the sum of two independent $\operatorname{Geom}(1/2)$ random variables. Thus, in the time-reversed process, points “un-merge” with interval $\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1$, where the geometrics are again independent. Thus, let $\rho^{(t)}$ be i.i.d. renewal processes with gap distribution
$Y\sim\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1,$
where the geometric random variables are independent and take the minimum value $1$. We have $\operatorname{\mathds{E}}Y=2+2-1=3$. The requisite coupling between the merged indices and the point processes $\rho^{(t)}$ used in the un-merging is obvious, and we omit it for the sake of brevity; see Figure 3 instead.
@@
Here, this summation corresponds to the fact that the same points contribute to two different gaps after one iteration of the time-reversed dynamics. The sequence $\overleftarrow{\eta}^{(t+1)}$ is indexed so that the 0–indexed point is the sum of the left-most element of the 0–th tuple and the right-most element of the (1)(-1)–st tuple. For example, in Figure 3, the un-mergings correspond to the entire gap that gets deleted in the merging, but also half of each of the preceding and following gaps.
@@ 3.4 Main Results
Moreover, the gap distribution of $\overrightarrow{\Xi}^{(\infty)}$ has an exponentially decaying tail.
We do not know the distribution of the Palm shift. The limit is not renewal, and thus it is difficult to identify; we leave it as an open problem. We use an appropriate martingale structure for the time-reversed process to establish Theorem 3.1. Using the standard Burkholder-Davis-Gundy inequality for discrete-time martingales gives an $O(p^{3/2})$ upper bound for the $p$-th moment, which is not enough to get exponential decay. To get an exponential tail decay, we directly control the MGF. The following corollary is a direct consequence of the proof technique of Theorem 3.1.
@@
The time-reversed process $\left({3}/{8}\right)^{t}\overleftarrow{\eta}^{(t)}$ is the stochastic dual of the gap sequence model, with respect to the inner product
$h(\overrightarrow{\Gamma}^{(s)},\left({3}/{8}\right)^{t}\overleftarrow{\eta}^{(t)}):=\left\langle\overrightarrow{\Gamma}^{(s)},\left({3}/{8}\right)^{t}\overleftarrow{\eta}^{(t)}\right\rangle=\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overrightarrow{\Gamma}^{(s)}_{i}\overleftarrow{\eta}^{(t)}_{i}$.
@@
Let $\overrightarrow{G}^{(t)}$ be the number of points of $\overrightarrow{\Xi}^{(t)}$ that have merged with the 0–indexed point of $\Xi^{(0)}$. Then
@@
for $2\leqslant p<\infty$. Moreover, the distribution of $\overrightarrow{G}^{(\infty)}$ has an exponentially decaying tail.
The following corollary follows by coupling the objects in the proofs of Theorems 3.1 and 3.3; but it is not a direct consequence of the theorem statements themselves.
@@
The joint weak convergence holds
$\left(\Theta\overrightarrow{\Xi}^{(t)},\left(\frac{3}{4}\right)^{t}\overrightarrow{G}^{(t)}\right)\Rightarrow\left(\Theta\overrightarrow{\Xi}^{(\infty)},\overrightarrow{G}^{(\infty)}\right).$
[…]
Define
$\overleftarrow{f}^{(t)}(x):=\overleftarrow{F}^{(t)}(x)-\overleftarrow{F}^{(t)}(x-(3/4)^{t})$
to be the mass of $\overleftarrow{F}^{(t)}(\cdot)$ at the point $x\in(3/4)^{t}\cdot\mathds{N}$.
Theorem 3.5.
The sequence of distribution functions $\overleftarrow{F}^{(t)}$ converges weakly almost surely to $\overleftarrow{F}^{(\infty)}$ as $t\to\infty$, where the (random) total mass of $\overleftarrow{F}^{(\infty)}$ is distributed as $\overrightarrow{\Xi}^{(\infty)}_{1}-\overrightarrow{\Xi}^{(\infty)}_{0}$, and the length of the (random) support of $\overleftarrow{F}^{(\infty)}$ is distributed as $\overrightarrow{G}^{(\infty)}_{1}$.
In Figure 4, we show simulations of $\overleftarrow{F}^{(t)}$ for two independent copies of our dynamics. We observe convergence as $t$ increases. However, we cannot directly depict the convergence of the measures via their distribution functions.
(a) First Simulation
(b) Second Simulation
We explicitly write out a dynamics for $\overleftarrow{F}^{(t)}$ in the following lemma that lies at the heart of creating a recursive structure for $\overleftarrow{F}^{(t)}$.
Lemma 3.6.
We have for all $t\geqslant 0,x\in\mathds{R}_{+}$,
$\overleftarrow{F}^{(t)}(x)+\frac{3}{4}\sum_{n\leqslant(4/3)^{t}x}\overleftarrow{f}^{(t)}((3/4)^{t}n)\cdot\bigl(\mathds{1}_{n\in\rho^{(t+1)}}-1/3\bigr)$
$\qquad=\overleftarrow{F}^{(t+1)}\bigl(x+(3/4)^{t+1}\cdot\bigl(|\rho^{(t+1)}\cap[0,(4/3)^{t}x]|-\frac{1}{3}(4/3)^{t}x\bigr)\bigr).$
For notational simplicity, we denote these dynamics as:
where
$\alpha^{(t+1)}_{x}$ $:=\frac{3}{4}\sum_{n\leqslant(4/3)^{t}x}\overleftarrow{f}^{(t)}((3/4)^{t}n)\cdot\bigl(\mathds{1}_{n\in\rho^{(t+1)}}-1/3\bigr),$
$\beta^{(t+1)}_{x}$ $:=(3/4)^{t+1}\cdot\bigl(|\rho^{(t+1)}\cap[0,(4/3)^{t}x]|-\frac{1}{3}(4/3)^{t}x\bigr).$
3.5 Concentration for Stationary Renewal Processes
Here, we present several technical lemmas and definitions used to prove our main results. The proofs of these lemmas can be found in Section 5.
Let $\rho$ be a stationary renewal process with gap distribution $Y\sim\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1$ and let $N^{\star}_{\rho}(x)=\left|\rho\cap[0,x]\right|,x\in\mathds{N}$ be the counting process. We have the following bound on the MGF of the excess count that will play a crucial role in the proof of Theorem 3.5. A more general result is given after this one.
Lemma 3.7.
For any $\theta_{0}>0$, there exists $A=A(\theta_{0})$ such that
$\operatorname{\mathds{E}}\left[e^{-\theta(x/\mu-N^{\star}_{\rho}(x))}\right]$ $\leqslant(1+A\left|\theta\right|)\cdot e^{A\theta^{2}x}$
$\text{ and }\operatorname{\mathds{E}}\left[e^{-\theta(x/\mu-N^{\star}_{\rho}(x))}\cdot\mathds{1}_{x\in\rho}\right]$ $\leqslant\frac{1}{\mu}(1+A\left|\theta\right|)\cdot e^{A\theta^{2}x}$
for all $x\in\mathds{N},\left|\theta\right|\leqslant\theta_{0}$.
Let $Y$ be a random variable having the interoccurrence time distribution with pmf $(f_{k})_{k\geqslant 1}$, mean $\mu=\operatorname{\mathds{E}}Y$ and mgf $\varphi(\theta):=\operatorname{\mathds{E}}e^{{\theta Y}}$ finite for some $\theta>0$. Let $Y_{*}$ denote a random variable with the equilibrium residual lifetime distribution with pmf
$\operatorname{\mathds{P}}(Y^{\star}=k)=f^{\star}_{k}=\frac{1}{\mu}\operatorname{\mathds{P}}(Y\geqslant k)\text{ for }k\geqslant 1.$
Let
$N_{t}$ $:=\max\{k\geqslant 0\mid 0<Y_{1}+Y_{2}+\cdots+Y_{k}\leqslant t\},$
$N^{\star}_{t}$ $:=\max\{k\geqslant 0\mid 0<Y^{\star}_{1}+Y_{2}+\cdots+Y_{k}\leqslant t\},\quad t=0,1,2,\ldots$
be the renewal counting process starting at 0 and at equilibrium, respectively. We denote by $\rho$ and $\rho^{\star}$ the corresponding point processes. We assume that for all $\alpha<0$,
[…]
For $\alpha\geqslant 0$, we define $C_{\alpha}=C^{\star}_{\alpha}=1$.
Lemma 3.8.
For $\theta\in\mathds{R}$, define the function $\psi(\theta):=\theta/\mu-\varphi^{-1}(e^{\theta})\geqslant 0$. Then, we have
$\operatorname{\mathds{E}}e^{-\theta(N^{\star}_{t}-t/\mu)}$ $\leqslant\max\left(\frac{1}{C^{\star}_{\alpha}},\frac{1}{C_{\alpha}}\right)\cdot\operatorname{\mathds{E}}e^{\alpha(Y^{\star}-1)}\cdot e^{t\psi(\theta)}$
$\text{ and }\operatorname{\mathds{E}}e^{-\theta(N^{\star}_{t}-t/\mu)}\mathds{1}_{{t\in\rho^{\star}}}$ $\leqslant\frac{1}{C_{\alpha}}\cdot\frac{1}{\mu}e^{t\psi(\theta)}$
for all $t\geqslant 0$ where $\alpha:=\varphi^{-1}(e^{\theta})$ has the same sign as $\theta$.
Remark 3.1.
Using the fact that $\psi(0)=\varphi^{-1}(1)=0,\psi^{\prime}(0)=1/\mu-(\varphi^{-1})^{\prime}(\varphi(0))=0,\psi^{\prime\prime}(0)=-(\varphi^{-1})^{\prime\prime}(1)=-1/\mu+\varphi^{\prime\prime}(0)\psi^{\prime}(0)/\mu^{2}=\operatorname{Var}(Y)/\mu^{3}$, for $\left|\theta\right|\ll 1$ we have $\psi(\theta)\approx{\operatorname{Var}(Y)}/{\mu^{3}}\cdot\theta^{2}/2.$
The following lemma will be useful for our example.
Lemma 3.9.
We have $\operatorname{\mathds{E}}e^{\alpha(Y^{\star}-1)}=\frac{\operatorname{\mathds{E}}e^{\alpha Y}-1}{\mu(e^{\alpha}-1)}$. Moreover, suppose that $Y$ satisfies the following condition: the function $k\mapsto\operatorname{\mathds{P}}(Y=k\mid Y\geqslant k)$ is increasing in $k$. Then $C_{\alpha}=e^{-\alpha}\operatorname{\mathds{E}}e^{\alpha Y}$ for all $\alpha<0$.
Finally, the following generalization of Lemma 3.8 will be used to derive exponential tail bounds for the limiting gap distribution.
Lemma 3.10.
For $t\geqslant 1$ and $\boldsymbol{v}=(v_{1},\dots,v_{t})\in\mathds{R}^{t}$, define $X_{t}:=\sum_{k=1}^{t}v_{k}(\mathds{1}_{k\in\rho^{\star}}-1/\mu).$ There exist constants $C,c\in(0,\infty)$, such that for all $t\geqslant 1$, all $\boldsymbol{v}\in\mathds{R}^{t}$, and all $\lambda\in\mathds{R}$ with $\left|\lambda\right|\cdot\left\|\boldsymbol{v}\right\|_{\infty}<c,$ we have
$\operatorname{\mathds{E}}e^{\lambda X_{t}}\leqslant e^{C\lambda^{2}\left\|\boldsymbol{v}\right\|_{2}^{2}}.$
3.6 Heuristics and a Toy Model
Here we consider a toy “follow-the-leader” model in which only averaging occurs without any folding, and it is easy to see the asymptotic behavior of the dual weight sequence. There are infinitely many agents located at positions given by $(X_{i}(0))_{i\in\mathds{Z}}$ on the real line where $X_{i}(0)<X_{i+1}(0)$ for all $i\in\mathds{Z}$. Over time $t\in\mathds{N}$, their locations $(X_{i}(t))_{i\in\mathds{Z}}$ change in the following way.
At discrete time $t\geqslant 1$, a Rademacher random variable $\varepsilon_{t}\sim\textrm{Uniform}\{-1,+1\}$ is chosen independently at random. If $\varepsilon_{t}=1$, all points move halfway towards their right neighbor; if $\varepsilon_{t}=-1$, all points move halfway towards their left neighbor to obtain their new positions at time $t$. Here, all points choose the same direction as the 0-indexed point, rather than taking independent decisions as in our model (Algorithm 1). In particular, no collisions can occur, and the ordering of the points is preserved at all times. No space scaling is required to keep the intensity constant.
Let $G_{i}(t)=X_{i}(t)-X_{i-1}(t),i\in\mathds{Z}$ be the gap sequence vector between consecutive neighbors at time $t$. Note that the evolution of the gap sequence at time $t$ can be written as an action of the random linear operator $\frac{1}{2}(T_{0}+T_{\varepsilon_{t}})$ applied to the previous gap sequence $\boldsymbol{G}(t-1):=(G_{i}(t-1))_{i\in\mathds{Z}}$ where $T_{a}$ is the linear shift by $a$ operator given by $(T_{a}\boldsymbol{x})_{i}=x_{i+a},\ i\in\mathds{Z}$. Moreover, the linear shift operators commute. Thus, we have
$\boldsymbol{G}(t)=\prod_{i=1}^{t}\frac{1}{2}(T_{0}+T_{\varepsilon_{i}})\cdot\boldsymbol{G}(0)=\frac{1}{2^{t}}\sum_{\sigma_{1},\sigma_{2},\ldots,\sigma_{t}\in\{-1,+1\}}T_{\varepsilon_{1}\sigma_{1}+\varepsilon_{2}\sigma_{2}+\cdots+\varepsilon_{t}\sigma_{t}}\cdot\boldsymbol{G}(0).$
Thus the gap sequence $G_{0}(t)$ can be written as the random linear combination $\sum_{k\in\mathds{Z}}a_{t}(k;\boldsymbol{\varepsilon}_{t})\cdot G_{k}(0)$ where $\boldsymbol{\varepsilon}_{t}=(\varepsilon_{1},\varepsilon_{2},\ldots,\varepsilon_{t})$ and the weights $a_{t}(i;\boldsymbol{v})$ as a function of $\boldsymbol{v}=(v_{1},v_{2},\ldots,v_{t})\in\{-1,+1\}^{t}$ is given by
$a_{t}(k;\boldsymbol{v})$ $=\operatorname{\mathds{P}}(v_{1}\xi_{1}+v_{2}\xi_{2}+\cdots+v_{t}\xi_{t}=-k),\qquad\xi_{1},\xi_{2},\ldots,\xi_{t}\stackrel{{\scriptstyle\textrm{i.i.d.}}}{{\sim}}\textrm{Uniform}\{0,1\}$
$=\operatorname{\mathds{P}}\biggl(\sum_{i=1}^{t}v_{i}\cdot(1-2\xi_{i})=2k+v_{1}+v_{2}+\cdots+v_{t}\biggr).$
Clearly, the weights are zero for $\left|k\right|>t$ and using the Central Limit Theorem we can easily see that the partial sum of the random weights $\sum_{i\leqslant k}a_{t}(i;\boldsymbol{\varepsilon}_{t})$ for $k\approx\theta\cdot\sqrt{t}-\frac{1}{2}\sum_{i=1}^{t}\varepsilon_{i},\ \theta\in\mathds{R}$ converges to Gaussian CDF $\Phi(2\theta)$. Thus, the total mass of the weights is constant here, and we have a “random” limiting distribution for the appropriately scaled partial sum of the weight sequence given by a Normal distribution with a random shift arising out of $t^{-1/2}\sum_{i=1}^{t}\varepsilon_{i}$.
@@ 4 Proofs of Main Results
Recall that $\overleftarrow{\eta}^{(t)}$ is a sequence of weights, from which we will construct (after a re-scaling by $\left(\frac{3}{8}\right)^{t}$) the gap sequence of the limiting point process. Note that $\overleftarrow{\eta}^{(t)}$ is supported on a subset of $\{-2^{t},-2^{t}+1,\ldots,2^{t}\}$, so that $\sum_{i=-\infty}^{\infty}{\overleftarrow{\eta}^{(t)}_{i}}$$\infty$. Define for $t\geqslant 0$ the process
@@
We can express the evolution of $\overleftarrow{M}^{(t)}$ as follows
$\overleftarrow{M}^{(t+1)}$ $=\left(\frac{3}{8}\right)^{t+1}\left(2\left\|\overleftarrow{\eta}^{(t)}\right\|_{1}+2\sum_{i\in\rho^{(t+1)}}\overleftarrow{\eta}^{(t)}_{i}\right)$
@@
$=\overleftarrow{M}^{(t)}+\frac{3}{4}\sum_{i\in\mathrm{supp}\overleftarrow{\eta}^{(t)}}\overleftarrow{X}^{(t)}_{i}\left(\mathds{1}_{i\in\rho^{(t+1)}}-\frac{1}{3}\right)=:\overleftarrow{M}^{(t)}+\overleftarrow{N}^{(t+1)}.$ (4.1)
Here, $\overleftarrow{X}^{(t)}_{\cdot}:=\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{\cdot}$. Notice that $\overleftarrow{N}^{(t+1)}$ has conditional mean 0. It follows that the process $\overleftarrow{M}^{(t)}$ is a unit-mean positive martingale; hence, it converges to an a.s. limit $\overleftarrow{M}^{(\infty)}$.
Notice that since $\max_{i\in\mathds{Z}}\overleftarrow{\eta}^{(t)}_{i}\leqslant 2^{t}$, we have that
$\max_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}\leqslant\left(\frac{3}{4}\right)^{t}.$
Equivalently, we have $\max_{i\in\mathds{Z}}\overleftarrow{X}_{i}^{(t)}\leqslant\left(\frac{3}{4}\right)^{t}$. Thus,
$\sum_{i\in\mathds{Z}}\left(\overleftarrow{X}_{i}^{(t)}\right)^{2}=\sum_{i\in\mathds{Z}}\left(\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}\right)^{2}\leqslant\left(\frac{3}{4}\right)^{t}\overleftarrow{M}^{(t)},$ (4.2)
@@
since $\overleftarrow{N}^{(t+1)}$ has conditional mean 0. Here, the $\sigma$-algebra $\mathcal{F}^{(t)}$ contains only those events related to the dynamics of $\overleftarrow{\eta}^{(t)}$: specifically, it does not contain events related to $\Gamma^{(0)}$. Substituting, we have
@@
$\leqslant\sum_{i\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\sum_{j\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\left(\frac{3}{4}\right)^{2t}\left|\operatorname{Cov}(\mathds{1}_{i\in\rho^{(t+1)}},\mathds{1}_{j\in\rho^{(t+1)}})\right|$
@@
Here, the convergence occurs using the fact that the renewal distribution of $\rho$ has an exponential tail and thus, $\operatorname{Cov}(\mathds{1}_{i\in\rho^{(t)}},\mathds{1}_{j\in\rho^{(t)}})$ decays exponentially in |ji||j-i|. To see this, let $Y\sim\mathrm{Geom}(1/2)+\mathrm{Geom}(1/2)-1$ denote the gap distribution of $\rho$ and let $f_{Y}(y):=\operatorname{\mathds{P}}(Y=y)$. Recall that $\mu:=\operatorname{\mathds{E}}Y=3$. Let $u_{0}=1$ and $u_{i}$ denote the probability of a renewal at time $i$ given that there is a renewal at time 0. Then $u_{n}=\sum_{i=0}^{n-1}u_{i}f_{Y}(n-i)$. We have $\operatorname{Cov}(\mathds{1}_{i\in\rho^{(t)}},\mathds{1}_{j\in\rho^{(t)}})=\frac{1}{\mu}(u_{|j-i|}-\frac{1}{\mu})$. From here, it is easily seen that $\sum_{i\geqslant 1}i\cdot\left|\operatorname{Cov}(\mathds{1}_{0\in\rho^{(t+1)}},\mathds{1}_{i\in\rho^{(t+1)}})\right|$ is summable. It follows that
@@
Figure 5: Relationship between the weight sequences $\eta^{(t)}$ and the initial gap sequence $\Gamma^{(0)}$. The orange dashed line, for example, is given by $\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{-1}+\overleftarrow{\Gamma}^{(0)}_{0}+\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{1}$; the blue dashed line is given by $\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{0}+\overleftarrow{\Gamma}^{(0)}_{1}+\overleftarrow{\Gamma}^{(0)}_{2}$.
Now, if the gap distribution of $\Xi_{0}$ has finite variance $\sigma^{2}$, we have the following (see Figure 5):
$\left\langle\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)},\overleftarrow{\Gamma}^{(0)}\right\rangle:=\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}\overleftarrow{\Gamma}^{(0)}_{i}$ $=\overleftarrow{M}^{(t)}+\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}(\overleftarrow{\Gamma}^{(0)}_{i}-1)$
$=:\overleftarrow{M}^{(t)}+\overleftarrow{L}^{(t)}.$
Obviously, we have $\operatorname{\mathds{E}}\overleftarrow{L}^{(t)}=0$ for all $t$, since $\operatorname{\mathds{E}}\Gamma^{(0)}_{i}=1$ for all $i\in\mathds{Z}$. Noting that $(\overleftarrow{\Gamma}_{i}^{(0)})_{i\in\mathds{Z}}$ are i.i.d. with variance $\sigma^{2}$, we have
$\operatorname{\mathds{E}}\left[\left(\overleftarrow{L}^{(t+1)}\right)^{2}\ \biggl|\ \mathcal{F}^{(t)}\right]$ $=\sum_{i\in\mathds{Z}}\sum_{j\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}_{i}^{(t)}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}_{j}^{(t)}\cdot\operatorname{Cov}(\overleftarrow{\Gamma}_{i}^{(0)},\overleftarrow{\Gamma}_{j}^{(0)})$